### v4

# Measurement-device agnostic quantum tomography

Abstract
Characterization of quantum states and devices is paramount to quantum science and technology. The characterization consists of individual measurements, which are required to be precisely known. A mismatch between actual and assumed constituent measurements limits the accuracy of this characterization. Here, we show that such a mismatch introduces reconstruction artifacts in quantum state tomography. We use these artifacts to detect and quantify the mismatch and gain information about the actual measurement operators. It consequently allows the mitigation of systematic errors in quantum measurement and state preparation.
I Introduction
The measurements play a central role in quantum science and technology [1]. However, the implemented measurements often differ from the ideal ones, leading to systematic errors. One can compensate for these systematic errors with the knowledge of the actual measurements. We investigate this problem from the point of view of quantum tomography, one of the most valuable tools in quantum science [2, 3, 4, 5, 6]. It allows the complete characterization of prepared states and implemented quantum circuits, assessing and certifying their quality. Quantum state tomography consists of a suitable set of known measurements on the investigated objects. One can reconstruct the state from the collection of measurement outcomes, the tomogram. Imperfect realization of these measurements causes a mismatch between the assumed and the realized measurement operators. Such a mismatch leads to the wrong interpretation of the acquired tomogram and manifests as a reconstruction artifact.
The true measurement operators can be fully characterized with measurement tomography [7, 8, 9, 10], which relies on perfect knowledge of input probe quantum states. Such a requirement is difficult to meet in the experiments. A self-consistent tomography framework has been introduced to loosen this condition [11]. The framework suggests the existence of a self-calibrating state class, which could reveal, when measured, the initially unknown information about the states and the measurement device. Another approach demonstrated for polarization states of light infers a single unknown parameter of the measurement device using maximum likelihood estimation [12]. There are also schemes that use simultaneous likelihood maximization over both states and devices [13] or states and measurements [14]. Finally, assuming the sparsity of the parameter vector and the low rank of the measured density matrix, the dependence on the measurement device can be relaxed [15].
In this work, we show that an ensemble of single-qubit states of equal purity distributed quasi-uniformly on the Bloch sphere forms a class of self-calibrating states. This class is convenient in experiments where states are naturally prepared with quasi-equal purity, e.g., in photonics and trapped ions. Advantageously, the precise control, characterization, or even sparsity of such self-calibrating probe states is not required. We show a novel approach to simultaneously calibrating quantum measurements and state preparation. To this end, we exploit artifacts in quantum state reconstruction performed on the self-calibrating probe states. In the multi-qubit setting, we can use the method in a fully scalable way to mitigate systematic errors on each local qubit measurement, effectively mitigating errors in the whole system. The method also applies to quantum device tomography thanks to the state-process duality [16, 17, 1], which allows the quantum processes to be treated as states. Our method is not limited to the determination of an initially unknown parameter shared among all measurement operators in the tomographic scheme; it allows the determination of parameters related to each measurement operator individually. The presented scheme provides the characterization of quantum states independently of measurement device imperfections. General concepts are introduced using numerical simulations, and we support our findings with a photonic experiment.
Figure 1: Purity modulation due to the measurement operator mismatch. (a-c) States are represented as points on the Bloch sphere plotted using Hammer projection, and the purity of their reconstruction is color-coded. Although all probe states were prepared with the same purity, the reconstruction using quantum tomography with mismatched measurement operators yields purity modulation. The purity modulation scales with the magnitude of error parameters. Panel (b) shows the purity modulation for randomly selected error parameters described in the main text. Panels (a) and (c) show how the purity modulation, if we scale the error vector by a factor of $\lambda=0.25$ and $\lambda=4$, respectively. The black curve in panel (d) depicts the scaling trend, and red dots on the curve mark the selected values of $\lambda$ plotted in panels (a-c).
II Measurement operator mismatch artifact
In quantum state tomography, many copies of an initially unknown state, described by density operator $\hat{\rho}$, are subject to a series of measurements described by measurement operators $\hat{\pi}_{j}$. These measurement operators should be tomographically complete, and their choice influences the performance of the tomography [18, 19, 20, 21]. The resulting collection of measurement outcomes forms the tomogram. The tomogram and the measurement operators are then input into an algorithm, which outputs the density operator. Here, we use the maximum-likelihood reconstruction algorithm [13, 22]. If the actual measurement operators $\hat{\pi}^{\prime}_{j}$ differ from those put into the reconstruction method, artifacts appear.
Let us demonstrate the behavior in the simplest scenario of pure probe single-qubit states and projective measurements onto the eigenstates of Pauli operators, called Pauli tomography. The measurement operator is realized by two rotation angles, corresponding to rotation about the $y$- and $z$-axis and subsequent projection onto computational basis state $|0\rangle\langle 0|$. Throughout the article, we use the computational basis $\{|0\rangle,|1\rangle\}$. The following operators describe the constituent measurements
$\hat{\pi}_{j}=\hat{R}_{z}^{\dagger}(\phi_{j})\hat{R}_{y}^{\dagger}(\theta_{j})|0\rangle\langle 0|\hat{R}_{y}(\theta_{j})\hat{R}_{z}(\phi_{j}),$ (1)
where $\hat{R}_{j}(\alpha)=\exp(i\hat{\sigma}_{j}\alpha/2)$ is the rotation generated by the Pauli operator. Then, the corresponding detection probability is given by the Born rule,
$p_{j}=\mathrm{Tr}[\hat{\rho}\hat{\pi}_{j}].$ (2)
The angles $\theta_{j}$ and $\phi_{j}$ for Pauli tomography are provided in Table S2 of the Supplemental Material. Assume that due to experimental imperfections, the actual angles are
$\theta_{j}^{\prime}=\theta_{j}+\delta_{j},$ (3)
$\phi_{j}^{\prime}=\phi_{j}+\epsilon_{j},$ (4)
i.e., we introduce systematic additive errors for each measurement operator. The error vector $\vec{\delta}=\{\delta_{j},\epsilon_{j}\}$ parameterizes the measurement operators $\hat{\pi}^{\prime}_{j}=\hat{\pi}_{j}(\vec{\delta})$. We numerically simulate tomograms using these true measurement operators for many pure probe states covering the Bloch sphere quasi-uniformly. Then, we reconstruct them assuming the original unperturbed measurement operators $\{\hat{\pi}_{j}\}$. This mismatch causes the state-dependent purity modulation, shown in Figure 1(a-c). This example shows the apparent purity modulation for the case of an error vector $\vec{\delta}$, whose elements are drawn from a normal distribution with $\sigma=5$ deg. We quantify the purity modulation in the ensemble of reconstructed states as
$\Delta P=\max P-\min P,$ (5)
where $P=\mathrm{Tr}\left(\hat{\rho}^{2}\right)$ is the purity of the reconstructed state.
To show how purity modulation scales with the error vector magnitude, we scale it with a non-negative real factor $\lambda$, i.e. $\vec{\delta}\rightarrow\lambda\vec{\delta}$, and observe the $\Delta P$, as shown in Figure 1(d). We see that $\Delta P$ monotonously increases with $\lambda$. Consequently, $\Delta P$ effectively measures how far the assumed measurement operators are from the actual ones. Moreover, the shape of the purity landscape is determined by the error vector $\vec{\delta}$. We will leverage this dependence of apparent purity as a function of vector $\delta$ to find the true measurement operators.
The method is not limited to the particular reconstruction method, as we illustrate in an example of rotating quarter-wave plate polarimetry with classical Fourier retrieval of Stokes parameters [23]. The example is provided in Supplementary Material S4. In that case, the disagreement between actual and assumed measurement operators causes inconsistency in the tomogram, which results in apparent state-dependent modulation of the degree of polarization, where the degree of polarization can even exceed one.
III Artifact mitigation
We minimize the artifact severity, i.e., purity modulation (5), by varying the measurement operators used to reconstruct the probe states. After recording all tomograms of the probe states we used, we optimize our assumption of $\vec{\delta}$ which specifies the measurement operators used in the reconstruction algorithm. Once the assumed parameters match the actual parameters, the artifact is minimized. Since the perturbed measurement operators no longer add the identity operator, we must properly renormalize them in the reconstruction [22].
Due to the state dependency, the Bloch sphere must be sampled densely enough to detect the mismatch. The probe states should cover the Bloch sphere quasi-uniformly and possess uniform purity; otherwise, their full control is not required. This feature is advantageous in practice because state preparation is never perfect in the experiment. It distinguishes our approach from the schemes that rely on perfect knowledge of the probe states, like quantum measurement tomography [8].
Generally, the optimization landscape has multiple local minima; therefore, global optimization must be performed. Here, we use the NOMAD implementation of the MADS algorithm [24]. The resulting optimum $\vec{\delta}$ is not unique because any set of measurement operators $\{\hat{U}\hat{\pi}_{j}\hat{U}^{\dagger}\}$, where $\hat{U}$ is a fixed unitary operation, also minimizes the artifact. These unitary-equivalent solutions could be interpreted as unitary transformations of probe states. Albeit this ambiguity, such a calibration is already applicable in schemes invariant to local unitaries such as random measurement [25] or entanglement certification [26, 27, 28], where only relative orientations of $\hat{\pi}_{j}$ operators are important. Moreover, the unitary operation can be easily determined by additional measurement of two known and non-orthogonal probe states.
Since the artifact is state-dependent, the probe sampling strategy is important for minimization. The sampling with too few probe states could lead to false error parameters $\vec{\delta}_{f}$. We found out that 30 quasi-uniformly distributed probe states provide a reliable determination of error parameters in all tested cases. When we used fewer probe states, we encountered incorrect estimates. The incorrect estimate provides reduced apparent purity modulation in the set of probe states, but when it is used on other states, the purity modulation can be even larger than initially.
To test the method numerically, we first randomly generate ground-truth error parameters $\vec{\delta}$. Each element is normally distributed with zero mean value and standard deviation of 10 deg. Then, we produce tomograms corresponding to pure probe states $\rho_{k}=|\psi_{k}\rangle\langle\psi_{k}|$. For simplicity, we exclude statistical noise from our simulations. Parameters $\vec{\delta}^{\prime}$ assumed in the reconstruction are optimized to minimize the purity modulation $\Delta P$ in the set of reconstructed probe states $\{\hat{\rho}_{k}\}$ for $k=1,\dots,30$, concluding the calibration step.
To verify the correctness of the optimization result, we use the ground truth $\vec{\delta}$ again to generate tomograms of pure test states $\hat{\xi}=|\xi_{k}\rangle\langle\xi_{k}|$, which quasi-uniformly covers the Bloch sphere in 108 points. We reconstruct these states using the optimized parameters $\vec{\delta}$ obtained in the calibration step and calculate the purity of the reconstructions. As a further test, we assume knowledge of two non-orthogonal probe states and use their reconstructions to determine corrective unitary operation $\hat{W}$. We apply $\hat{W}$ to all reconstructed test states $\hat{\xi}_{k}$ to eliminate the unitary-related ambiguity. Then, we check their fidelity with the expected ideal test state, $F_{k}=\mathrm{Tr}[\sqrt{\sqrt{\hat{\xi}_{k}}\hat{\rho}_{k}\sqrt{\hat{\xi}_{k}}}]$. We characterize the test states ensemble with the purity modulation and their lowest fidelity $\min\limits_{k}F_{k}$ to provide the worst-case estimate.
We repeated the whole numerical simulation 100 times, randomly choosing new ground truth for the error parameters $\delta$ in each run. The details about optimizer settings are provided in Section S1 of the Supplementary Material. When we compare the optimized results to the reference, i.e. the states reconstructed with the assumption of null error vector, the average purity modulation decreases from $\Delta P=(8\begin{subarray}{c}+3\\ -3\end{subarray}))\cdot 10^{-2}$ to $\Delta P=(1.0\begin{subarray}{c}+0.8\\ -1.0\end{subarray})\cdot 10^{-3}$. The uncertainty interval spans from 0.158 to 0.842-quantile in correspondence to one standard deviation and is used due to the skewed distribution of the results. Mainly, the infidelity decreased from $(3\begin{subarray}{c}+1\\ -2\end{subarray})\cdot 10^{-2}$ to $(3\begin{subarray}{c}+5\\ -3\end{subarray})\cdot 10^{-4}$, which represents two orders of magnitude improvement in accuracy. This improvement is crucial for reaching error levels low enough for quantum error correction schemes. The lower $\Delta P$ of the probe state ensemble was achieved the better results in the test ensemble we achieved. These results illustrate the feasibility of such optimization and its ability to reliably determine the true measurement operators. It is important to stress that the pure probe states and projective (pure) measurements were chosen here for the simplicity of introducing the measurement-device agnostic framework. However, the method is fully applicable in real cases of generic positive-operator-valued measurements and nonunity purity, as indicated in the following experimental demonstration.
In the multi-qubit scenario, the tomography is usually done using local measurements. In that case, we use the presented method for each local measurement device. The artifact severity scales worst as $P_{\min,n}=\prod\limits_{i=1}^{n}P_{\min,i}$ for $n$-qubit state tomography, where $P_{\min,i}$ is the corresponding minimal purity observed at the single qubit. If the input state |$\ket{\psi_{i}}$ at the $i$-th qubit suffers from the artifact the most, then product state $\bigotimes\limits_{i}\ket{\psi_{i}}$ is the $n$-qubit state with the greatest manifestation of the artifact. In such a case, the single-qubit improvements multiply and lead to significantly better overall results in multi-qubit scenarios.
IV State preparation
The state preparation typically suffers from imperfections too. Here, we use the symmetry in the Born rule to formally exchange state preparation and measurement to utilize the previously introduced method to correct state preparation. Usually, the qubit is initialized in state $|0\rangle$, then turned into state $|\psi_{j}\rangle=\hat{U}_{j}|0\rangle$ by unitary evolution $\hat{U}_{j}$. The tomographic projection is also typically realized by some unitary evolution $\hat{V}_{k}$ and subsequent projection, i.e., $|\pi_{k}\rangle=\hat{V}_{k}^{\dagger}|0\rangle$. The state tomogram of the probe state $|\psi_{j}\rangle$ consists of measured probabilities $|\langle\hat{\pi}_{k}|\psi_{j}\rangle|^{2}=|\langle 0|\hat{V}_{k}\hat{U}_{j}|0\rangle|^{2}$. The discussed procedure obtains information related to projections, i.e. about $V_{k}$.
Physical reversal of the process consists of preparing states $\hat{V}_{k}|0\rangle$ and projecting them onto states $\hat{U}_{j}^{\dagger}|0\rangle$, obtaining tomogram elements $p_{kj}\propto|\langle 0|\hat{U}_{j}\hat{V}_{k}|0\rangle|^{2}$. We can analyze the tomograms to improve our knowledge of the measurement operator, i.e., $\hat{U}_{j}^{\dagger}|0\rangle\langle 0|\hat{U}_{j}$, related to state preparation in the non-reversed situation. The physical reversal can be realized in the experiment by performing projections onto the original probe states in the tomographic device and preparing states corresponding to the original tomographical projectors instead of the original probes. Either way, we gain information about $\hat{U}_{j}$, increasing the accuracy of state preparation.
V Special case: multiplicative error
An important subclass of the problem is the case of multiplicative measurement operator error, where
$\theta_{j}^{\prime}=(1+\delta)\theta_{j},$ (6)
$\phi_{j}^{\prime}=(1+\epsilon)\phi_{j}.$ (7)
This case is relevant when we sequentially switch the unitary operations that precede a fixed projector to perform the tomographic measurements. These multiplicative errors could be interpreted as Bloch sphere under- or over-rotation. This approach to quantum state tomography has been reported in various platforms, including optics [29, 30], color-centers in solids [31], trapped ions [32], neutral atoms [33], or superconducting qubits [34], making this problem subclass highly relevant.
Because parameters $(\delta,\epsilon)$ are shared among all projection operators, the optimum search simplifies from 12 parameters to just two. In Section S2 of the Supplementary material, we show that with just eight probe states, the optimization landscape possesses a single minimum, and therefore standard gradient-descent optimization is applicable. The advantage is better performance when compared to global optimization and also the smaller probe state set. Another advantage is that the global optimum is now unique, and the method unambiguously finds the true set $\{\hat{\pi}_{j}\}$ even when there is an unknown unitary operation between preparation and tomographic measurements.
Figure 2: Experimental demonstration for a polarization-encoded photonic qubit. Panel (a) shows the experimental scheme: LD - laser diode, HWP (QWP) - half-(quarter)-wave plate, GT - Glan-Taylor polarizer, BD - calcite beam displacer, PD - photodiode. Panels (b) and (c) show the purity and fidelity of reconstructed test states before (blue) and after (orange) application of measurement-device agnostic calibration. The fidelity of reconstructed states is calculated with respect to theoretical test states. The theoretical test states are depicted as empty black circles on the projected Bloch sphere in panels (d, e). Panel (d) depicts the reference case – the method was not used, and we assumed perfect waveplates in preparation and tomographic projections. In contrast, panel (e) shows the improved precision of preparation and measurement when we use the method and take into account the retardance deviation. The purity modulation is reduced and the test states match closely the theoretical expectations. In both panels, the reconstructed states are represented on the Bloch sphere using Hammer projection, and their purity is color-coded.
VI Experiment: polarization state tomography
We experimentally demonstrate the method in the waveplate-based polarization state tomography scenario, an example of a situation with multiplicative errors. At the end of the process, we will improve our knowledge of the waveplate retardance and consequently mitigate the reconstruction artifact.
The experimental setup is depicted in Fig. 1(a). We used a continuous-wave 810 nm thermally-stabilized fiber-coupled laser diode (LD). The laser beam was decoupled into free space using an 11-mm aspheric lens and horizontally polarized using a Glan-Taylor prism (GT). The waveplates HWP1 and QWP1 effectively control the prepared polarization state. The following operator describes the action of a waveplate:
$\hat{W}_{\alpha,\gamma}=|\alpha\rangle\langle\alpha|+e^{-i\Gamma}|\alpha_{\bot}\rangle\langle\alpha_{\bot}|,$ (8)
where $\alpha$ and $\Gamma$ are its angular position and retardance, and $|\alpha\rangle=\cos{\alpha}|0\rangle+\sin{\alpha}|1\rangle$ is a linearly polarized state, and $\alpha_{\bot}=\alpha+\pi/2$. To prepare a state located at colatitude $\theta$ and longitude $\phi$ on the Bloch sphere, we rotate the half- and quarter-wave plate to angles $x^{\prime}$ and $y^{\prime}$ given by the relations
$y^{\prime}=-\frac{1}{2}\arcsin\left(\sin\theta\sin\phi\right),$ (9)
$x^{\prime}=\frac{1}{4}\arctan\left(\tan\theta\cos\phi\right)+\frac{1}{2}y+c(\theta),$ (10)
where $c(\theta)=\pi/4$ if $\theta>\pi/2$ and $c(\theta)=0$ otherwise. This relation holds for wave plates that have perfect retardance. The deviation from the ideal retardance will introduce errors in the preparation. We assumed the perfect retardance in the first step of the experiment as our method does not require exact knowledge of probe states.
To demonstrate the tolerance of the method to a unitary operation between preparation and tomographic measurement, we add a quarter-wave plate (QWPX) with its fast axis oriented at -12.5 deg relative to the horizontal direction. Because this unitary operation can be in principle unknown, we restrict ourselves from using any information about QWPX in the calibration.
The states were analyzed using half-waveplate and quarter-wave (HWP2, QWP2) and subsequent projection onto horizontal and vertical polarization using a calcite beam-displacer (BD). The waveplate settings for the Pauli tomography are listed in Tab. S2 of the Supplementary Material. All waveplates were mounted into a precision motorized rotation stage that provided bidirectional repeatability of ±$\pm 30$ mdeg.
We used reverse-biased PIN photodiodes (PD) to measure optical intensity at both outputs of the beam-displacer. We normalized the readings using the sum of two signals to eliminate laser power fluctuation.
The action of the waveplate could be seen as the rotation of the Bloch sphere around axis $(\sin(2\alpha),0,\cos(2\alpha))$ with rotation angle $\Gamma$, where $\alpha$ determines the angular position of the waveplate and $\Gamma$ its retardance. The retardance usually slightly deviates from its nominal value. These deviations manifest as under- or over-rotation of the Bloch sphere. We parameterize the measurement operator by precisely controlled waveplate angular positions $x$, $y$, and half- and quarter-wave plate retardance deviations $\delta$ and $\epsilon$, respectively. These deviations are initially unknown to us, and we will use the presented method to estimate their values.
We use eight probe states and perform the Pauli tomography for each. Initially, we assume zero deviations, $\delta=\epsilon=0$, and observe the purity modulation $\Delta P=0.0572(3)$. The number in parenthesis is one standard deviation at the last significant digit, e.g. $0.0572(3)=(572\pm 3)\cdot 10^{-4}$. The uncertainties were estimated using bootstrapping with the assumption of normally distributed photo-current with 0.1% relative standard deviation, which is the most dominant term and is related to the ammeter precision. The shot noise is negligible at the power level of our laser, compared to the technical noise. Then, we vary parameters $\delta$ and $\epsilon$ and update the measurement operators employed in the reconstruction to minimize the purity modulation. The optimal $\delta$ and $\epsilon$ are close to the true values because the reconstruction artifact is significantly reduced to $\Delta P=0.0140(2)$. We found the optimal parameters $\delta=4.5(2)$ deg, $\epsilon=-1.30(4)$ deg.
We then reverse the process to determine retardance deviations $\tilde{\delta},\tilde{\epsilon}$ of the waveplates used in state preparation. We prepare the six eigenstates of Pauli operators, project each onto eight states corresponding to the original probe state set, and treat this data as before. We minimize the apparent purity modulation from 0.0700(3) to 0.0144(2) to find the optimal parameters $\tilde{\delta}=4.53(6)$ deg and $\tilde{\epsilon}=-3.58(3)$ deg.
The calibration step revealed the retardance deviation of the wave plates used for preparation and projection. To reconstruct the quantum states, we could directly use this knowledge to update our measurement operators. However, we take one extra step which allows us to prepare the states with greater precision and control the measurement operators precisely. We took these deviations into account in the verification step and calculated new angular positions of wave plates for tomographic projections and state preparation. Since Eq. 10 does not hold anymore, we used numerical optimization to find the desired wave plate settings ($x,y$):
$x,\,y=\mathrm{argmax}_{x,y}|\langle\psi|\hat{W}(x,\pi+\delta)\hat{W}(y,\pi/2+\epsilon)|0\rangle|^{2}.$ (11)
Similarly, to find the settings to project a state onto state $|\pi\rangle$ we perform this maximization
$x,\,y=\mathrm{argmax}_{x,y}|\langle\langle 0|\hat{W}(x,\pi/2+\epsilon)\hat{W}(y,\pi+\delta)|\pi\rangle|^{2}.$ (12)
At this point, we have information about actual measurement operators, and we have also calibrated the state preparation. This calibration was done in the presence of a unitary operation between state preparation and measurement, which was introduced by QWPX. This operation was effectively unknown because we did not use any prior information about QWPX during the process.
Now, we proceed to certify the correctness of such a calibration. The calibration should remain valid even if we change the central unitary operation introduced by QWPX. To test this, we removed the QWPX and then prepared an ensemble of 58 test states, depicted as empty circles in Figure 2(c,d).
As a reference, we first assumed perfect waveplates and performed state tomography for each test state. We indeed observed purity modulation $\Delta P=0.0607(2)$ with purity minimum $P_{\min}=0.9383(2)$. This modulation indicates reconstruction artifacts. The minimum fidelity of the reconstructed state to the target states was $F_{\min}=0.96688(4)$. These fidelity values indicate imprecise preparation.
In contrast, when we used our knowledge of $\delta$ and $\epsilon$ in the reconstruction and preparation, the minimum purity increased to $0.9830(3)$ and the minimum fidelity to $F_{\min}=0.9907(2)$. The information gained in the calibration clearly improved the reconstruction.
The results in Figure 2 indicate improved quality of state preparation and tomographic measurements when we use the knowledge of $\delta$ and $\epsilon$ to improve both preparation and measurement. The small remaining artifacts are present due to the limited precision of the learned retardance deviations and also the repeatability of the waveplate rotation.
VII Path-encoded qubits on a photonic chip
Figure 3: Application of the method for on-chip photonic path qubit. (a) Projection apparatus for a path qubit. Applied voltage $V_{1,2}$ on heater elements (orange) causes thermal-based phase shift in the waveguide interferometers. These shifts determine the measurement projector. The photons in the upper output path are detected with a single-photon detector (SPAD). (b) 1D-section of 6D function $\Delta P(c_{1},\dots,c_{6})$. The value of $\Delta P$ at the initial assumption of $c_{j}$ is $1.8\cdot 10^{-2}$ on the ensemble of probe states. Minimum $\Delta P=4.5\cdot 10^{-4}$ lies at the true value $c^{\prime}_{6}$. (c) Initial mismatch results in state-dependent purity artifact. We plot test states on the Bloch sphere as circles and color-code their reconstructed purity. (d) When we consider the true retardance, the artifact vanishes.
As another example, we numerically illustrate the measurement-device agnostic calibration for tomography of photonic on-chip path qubits [30, 35, 36, 37]. The qubit is encoded into the photon propagation in one or another waveguide (states $\ket{\uparrow}$ or $\ket{\downarrow}$), as depicted in Figure 3(a). The local heating of one waveguide controls the phase between the two paths. The heating is typically achieved by an electrical current flowing through a small heater element. A Mach-Zehnder interferometer serves as a variable coupler, and its interferometric phase controls the coupling ratio. Detection of photons in a single optical path implements a projective measurement of the path qubit. The circuit in Figure 3(a) implements a projection onto state |$\ket{\pi}$, whose colatitude and longitude on the Bloch sphere are controlled by optical phases. The relation between optical phases $\varphi_{1,2}$ and voltages $V_{1,2}$ applied on the heating elements is approximated with polynomials [35]
$\varphi_{1}(V_{1})=c_{1}+c_{2}V_{1}+c_{3}V_{1}^{2},$ (13)
$\varphi_{2}(V_{2})=c_{4}+c_{5}V_{2}+c_{6}V_{2}^{2}.$ (14)
Here, the voltages $V_{1,2}$ are controlled perfectly, but parameters $c_{j}$ are known only approximately. The other imperfections are neglected. We have increased the number of probe states to 22 to compensate for the higher number of optimization parameters. The simulated experiment, depicted in Fig. 3(c), shows the measurement operator mismatch artifact. We search the six-dimensional parametric space to find the point of minimal $\Delta P(c_{j})$ where the systematic error is mitigated. Panel (b) shows a one-dimensional section illustrating that the minimal $\Delta P$ corresponds to the true value $c^{\prime}_{6}$. Taking the true values $c^{\prime}_{j}$ into account mitigates the reconstruction artifact, as shown in panel (d).
VIII Conclusion
The mismatch between actual measurement operators and their theoretical counterparts assumed in reconstructing quantum states or devices introduces reconstruction artifacts. We showed how to leverage these artifacts to reveal the actual measurement operators and mitigate systematic errors in the tomography and state preparation. We experimentally applied this method to Pauli tomography of polarization-encoded photonic qubits and numerically analyzed several other measurement schemes. The proposed method makes the tomography independent of the measurement device. The main advantage is that the method does not require perfect knowledge or control of the probe states, which is virtually impossible to obtain. The method can calibrate already existing tomographic apparatus without the need for individual characterization of its constituent components. This is particularly useful in the case where the experimental setup is monolithic and individual components cannot be calibrated individually, e.g. in the case of integrated circuits. One can also calibrate the state preparation using the same method. Furthermore, the presented approach applies locally to individual parties of a larger multiparty system; i.e. it is fully scalable.
In summary, we developed the framework for accurate and scalable calibration of quantum measurement devices and also quantum state preparation, which is critical in any applications of quantum science and technology.
The application of the reported measurement-device agnostic quantum measurement goes beyond the full quantum tomography [13, 22]. The same method will find its use for mitigation of measurement errors in various approximative and more scalable approaches such as Monte-Carlo sampling [38, 39, 40, 41], permutationally invariant tomography [42], compressed sensing [43, 44], etc. The method can also improve the accuracy of building and characterization of the photonic and quantum circuits in complex media [45, 46]. Finally, the reported approach also applies to experimental realizations of entanglement certification and quantification [26, 27, 25, 47, 28].
The supporting data for this article are available from the Zenodo repository [48].
Acknowledgements.
We acknowledge the support of the Czech Science Foundation under grant No. 21-18545S. M. B. acknowledges the support of Palacký University under grant No. IGA-PrF-2024-008. We acknowledge the use of cluster computing resources provided by the Department of Optics, Palacký University Olomouc. We thank J. Provazník for maintaining the cluster and providing support.